# Contraherent cosheaves of contramodules on Noetherian formal schemes

### 29 Mar 2026 v1

@@ Abstract.
We define the exact category of contraherent cosheaves of contramodules on a locally Noetherian formal scheme, as well as the exact categories of locally contraherent cosheaves of contramodules (with respect to a given open covering). The exposition in the section of preliminaries in adic commutative algebra is worked out in the greater generality of arbitrary commutative rings with adic topologies (of finitely generated ideals).
@@ 1. Introduction
One of the main original goals of the theory of contraherent cosheaves, developed in [17, 29] and discussed in [28], was to globalize contramodules over formal schemes and ind-schemes. In the present paper, we accomplish this goal in the modest generality of (locally) Noetherian formal schemes with adic topologies on the rings of functions on affine open formal subschemes. These are the formal schemes in the sense of Hartshorne’s textbook [12, Section II.9], as distinguished from the more general approaches of Grothendieck’s EGAI [11, Sections 0.7 and I.10] and The Stacks project [13, Chapter Tag 0AHW].
Why contraherent cosheaves of contramodules? Why not quasi-coherent sheaves of contramodules? This question was raised in [28, Section 2] and answered in [32, Section 1.10], with a definitive counterexample in [32, Remark 10.7 and Example 10.8]. Simply put, the localization functors in contramodule categories do not have the requisite exactness properties (which they have in the usual module categories). The exactness properties of the colocalization functors in the module categories are better. Another explanation is that one wants to work with infinite direct products of contramodules, and the localization functors do not commute with direct products. The main results of the paper [32] establish the existence, an explicit construction, and the exactness properties of the colocalization functors in contramodule categories in a generality far surpassing the needs of the present paper.
@@ 2.1. Adic topological rings and their maps
Let $R$ be an adic topological ring and $I\subset R$ be an ideal of definition of $R$. Then, for every $n\geq 1$, the ideal $I^{n}\subset R$ is an ideal of definition, too. Another given ideal $J\subset R$ is an ideal of definition if and only if there exist integers $n$, $m\geq 1$ such that $I^{n}\subset J$ and $J^{m}\subset I$. For any two ideals of definition $I$, $J\subset R$, the ideals $I+J$,  $I\cap J$, and $IJ\subset R$ are also ideals of definition in $R$.
@@
We will say that an adic topological ring $R$ is adically Noetherian (cf. [31, Section 9]) if it satisfies the equivalent conditions of Lemma 2.1.2.
@@
Let $R$ and $S$ be two commutative rings with adic topologies, and let $f\colon R\longrightarrow S$ be a ring homomorphism. Then the following three conditions are equivalent:
@@
    for every ideal of definition $I\subset R$ there exists an ideal of definition $J\subset S$ such that $J\subset Sf(I)$.
@@
(3) $\Longrightarrow$ (4) $\Longrightarrow$ (1) Obvious. ∎
@@ 2.2. Torsion modules
Let $R$ be a commutative ring and $s\in R$ be an element. Following the notation in [17, Section 1.1], [20, Section 2.1], etc., we denote by $R[s^{-1}]$ the ring $R$ with the element $s$ formally inverted. In other words, $R[s^{-1}]=S^{-1}R$ is the localization of $R$ at the multiplicative subset $S=\{1,s,s^{2},s^{3},\dotsc\}\subset R$.
@@
Let $I\subset R$ be an ideal. We say that an $R$-module $\mathcal{M}$ is $I$-torsion if $\mathcal{M}$ is $s$-torsion for all $s\in R$. One can easily see that it suffices to check this condition for the element $s$ ranging over any chosen set of generators of the ideal II[30, Lemma 1.2].
Other definitions of the term “$I$-torsion $R$-module” exist in the literature (see, e. g., [14, Section 3]), but the difference between them does not manifest itself for finitely generated ideals $I$, which we are really interested in. So, for a finitely generated ideal $I\subset R$, an $R$-module $\mathcal{M}$ is $I$-torsion if and only if, for every element $x\in\mathcal{M}$, there exists an integer $n\geq 1$ such that $I^{n}x=0$ in $\mathcal{M}$[30, Lemma 1.3].
@@
The assertions is obvious; see [20, Lemma 6.1(a)] for some details. ∎
@@
An $R$-module $\mathcal{M}$ is torsion if and only if, for every element $x\in\mathcal{M}$, the annihilator of $x$ in $R$ is an open ideal. In other words, an $R$-module $\mathcal{M}$ is torsion if and only if it is discrete (as a module over the topological ring $R$) in the sense of [39, Section VI.4], [3, Section 1.4], [15, Section D.5.1], [17, Section E.2], [18, Sections 1.4 and 2.1], [22, Section 2.3], [23, Section 2.4], [35, Section 7.2], [36, Section 1], etc.
@@
For any ideal $I$ in a commutative ring $R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ is closed under submodules, quotients, extensions, and infinite direct sums in $R{\operatorname{\mathsf{--Mod}}}$. In other words, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ is a localizing subcategory, or in a different terminology, a hereditary torsion class [39, Sections VI.2–3] in $R{\operatorname{\mathsf{--Mod}}}$.
@@ 2.3. Adic completions
There is a natural completion map $\lambda_{I,M}\colon M\longrightarrow\Lambda_{I}(M)$. We say that an $R$-module $M$ is $I$-adically separated if the map $\lambda_{I,M}$ is injective, and that $M$ is $I$-adically complete if $\lambda_{I,M}$ is surjective. For any $R$-module $M$, the $R$-module $\Lambda_{I}(M)$ is $I$-adically separated and complete [40, Corollary 3.6]. The latter assertion depends on the assumption that the ideal $I\subset R$ is finitely generated [40, Example 1.8]. The functor $\Lambda_{I}$ is left adjoint to the inclusion of the full subcategory of $I$-adically separated and complete $R$-modules into $R{\operatorname{\mathsf{--Mod}}}$, and the map $\lambda_{I,M}$ is the adjunction unit [20, Theorem 5.8].
@@
The converse assertion to Lemma 2.3.1(a) is not true: There exist continuous maps of adic topological rings $f\colon R\longrightarrow S$ such that the continuous ring map $\Lambda(f)$ is tight but the map $f$ is not. For a counterexample, it suffices to take $R=S=k[x]$ and $f=\mathrm{id}_{k[x]}$ as in Examples 2.1.7, with the discrete ($(0)$-adic) topology on $R$ and the indiscrete ($(1)$-adic) topology on $S$. Then one has $\Lambda(R)=R$ and $\Lambda(S)=0$; so the ring map $\mathfrak{f}=\Lambda(f)$ is tight; but the ring map $f$ is not.
Let $R$ be an adic topological ring and $\mathfrak{R}=\Lambda(R)$ be its completion. The adic completion functor $\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ or $\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ is neither left, nor right exact [40, Example 3.20]. This fact is closely related to the fact that the category of separated and complete $R$-modules is not abelian [20, Example 2.7(1)]. To overcome these difficulties, one considers contramodule $R$-modules instead of the separated and complete $R$-modules; see the next section.
@@ 2.4. Contramodules
Let $R$ be a commutative ring and $s\in R$ be an element. An $R$-module $P$ is said to be an $s$-contramodule if $\operatorname{Hom}_{R}(R[s^{-1}],P)=0=\operatorname{Ext}_{R}^{1}(R[s^{-1}],P)$. The ring $R[s^{-1}]$ is considered as an $R$-module here. The definition above does not mention any $\operatorname{Ext}^{i}$ with $i\geq 2$ because the projective dimension of the $R$-module $R[s^{-1}]$ never exceeds 11[17, Section 1.1], [16, proof of Lemma B.7.1], [20, proof of Lemma 2.1]. A detailed discussion of this definition, which goes back to [15, Remark A.1.1], [16, Theorem B.1.1], [17, Section D.1], and [19, Section 2], can be found in [20, Sections 1–3].
Let $I\subset R$ be an ideal. An $R$-module $\mathfrak{P}$ is said to be an $I$-contramodule (or an $I$-contramodule $R$-module) if $\mathfrak{P}$ is an $s$-contramodule for every $s\in I$. It suffices to check this condition for the element $s$ ranging over any chosen set of generators of the ideal II[16, Theorem B.1.1], [19, Section 2]. A detailed discussion with two proofs can be found in [20, Theorem 5.1].
@@
This is [20, Lemma 6.1(b)]. ∎
Assume that the ideal $I\subset R$ is finitely generated. Then any $I$-contramodule $R$-module is $I$-adically complete [20, Theorem 5.6], but it need not be $I$-adically separated [37, Example 2.5], [40, Example 3.20], [14, Example 4.33], [15, Section A.1.1], [18, Section 1.5], [20, Example 2.7(1)]. Any $I$-adically separated and complete $R$-module is an $I$-contramodule [20, Lemma 5.7].
An $I$-contramodule $R$-module is said to be quotseparated [34, Section 5.5], [24, Section 1] if it is a quotient $R$-module of an $I$-adically separated and complete $R$-module. The references in the previous paragraph provide examples of quotseparated $I$-contramodule $R$-modules that are not $I$-adically separated. For an example of an $I$-contramodule $R$-module that is not quotseparated, see [19, Example 2.6], [24, Examples 1.8]. Over a Noetherian commutative ring $R$, all $I$-contramodule $R$-modules are quotseparated [16, Theorem B.1.1], [24, Corollary 3.7 and Remark 3.8]. A perhaps more detailed discussion can be found in [31, Section 1].
Now let $R$ be a commutative ring endowed with an adic topology in the sense of Section 2.1. Let $I$, $J\subset R$ be two ideals of definition of the ring $R$. Then an $R$-module $\mathfrak{P}$ is an $I$-contramodule if and only if it is a $J$-contramodule [20, Remark 5.5]. If this is the case, we will say that $\mathfrak{P}$ is a contramodule $R$-module (with respect to the given adic topology of $R$). Moreover, a contramodule $R$-module is quotseparated as an $I$-contramodule $R$-module if and only if it is quotseparated as a $J$-contramodule $R$-module; so one can speak of quotseparated contramodule $R$-modules.
@@
For any ideal $I$ in a commutative ring $R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is closed under kernels, cokernels, extensions, and infinite direct products in $R{\operatorname{\mathsf{--Mod}}}$[7, Proposition 1.1], [19, Section 2], [20, Theorem 1.2(a)]. It follows that $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is an abelian category. In fact, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is a locally $\aleph_{1}$-presentable abelian category with enough projective objects [33, Example 4.1(3)] (see also [21, Example 1.3(4)]). The identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is exact and preserves infinite products (hence also all limits) and $\aleph_{1}$-directed colimits, but not infinite direct sums. The infinite direct sum functors in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ are usually not exact.
@@
In the case of a finitely generated ideal $I\subset R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ is closed under kernels, cokernels, and infinite direct products in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ and in $R{\operatorname{\mathsf{--Mod}}}$, and also under subobjects and quotient objects in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$[24, Lemma 1.2]. (However, the full subcategory of quotseparated contramodules is not closed under extensions; in fact, any $I$-contramodule $R$-module is an extension of two quotseparated $I$-contramodule $R$-modules [24, Proposition 1.6].) It follows that $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ is an abelian category. In fact, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is also a locally $\aleph_{1}$-presentable abelian category with enough projective objects [24, end of Section 1]. The identity inclusion functors $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ and $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ are exact and preserve infinite products (hence also all limits) and $\aleph_{1}$-directed colimits. The inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ does not preserve infinite direct sums. The infinite direct sum functors in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ are usually not exact (the global dimension $1$ case is one major exception [16, Remark 1.2.1]).
@@
To any complete, separated topological ring $\mathfrak{R}$ where open right ideals form a base of neighborhoods of zero, the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ of left $\mathfrak{R}$-contramodules is naturally assigned [15, Remark A.3], [16, Section 1.2], [18, Section 2.1], [33, Sections 1.2 and 5], [22, Section 2.7], [23, Sections 2.6–2.7], [35, Section 6], [36, Section 1]. The category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ comes equipped with a natural exact, faithful forgetful functor $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$, preserving infinite products (hence all limits).
@@
In the case of an adic topological ring $R$, the composition of forgetful functors $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful, and its essential image coincides with the full subcategory of quotseparated contramodule $R$-modules [24, Proposition 1.5]. So there is a natural equivalence (in fact, isomorphism) of categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\simeq R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$.

### 17 Apr 2026 v2

@@ Abstract.
We define the exact category of contraherent cosheaves of contramodules on a locally Noetherian formal scheme, as well as the exact categories of locally contraherent cosheaves of contramodules (with respect to a given open covering). We also construct the direct image and inverse image functors of locally contraherent cosheaves of contramodules under morphisms of locally Noetherian formal schemes. The exposition in the section of preliminaries in adic commutative algebra is worked out in the greater generality of arbitrary commutative rings with adic topologies (of finitely generated ideals).
@@ 1. Introduction
One of the main original goals of the theory of contraherent cosheaves, developed in [18, 30] and discussed in [29], was to globalize contramodules over formal schemes and ind-schemes. In the present paper, we accomplish this goal in the modest generality of (locally) Noetherian formal schemes with adic topologies on the rings of functions on affine open formal subschemes. These are the formal schemes in the sense of Hartshorne’s textbook [13, Section II.9], as distinguished from the more general approaches of Grothendieck’s EGAI [12, Sections 0.7 and I.10] and The Stacks project [14, Chapter Tag 0AHW].
Why contraherent cosheaves of contramodules? Why not quasi-coherent sheaves of contramodules? This question was raised in [29, Section 2] and answered in [33, Section 1.10], with a definitive counterexample in [33, Remark 10.7 and Example 10.8]. Simply put, the localization functors in contramodule categories do not have the requisite exactness properties (which they have in the usual module categories). The exactness properties of the colocalization functors in the module categories are better. Another explanation is that one wants to work with infinite direct products of contramodules, and the localization functors do not commute with direct products. The main results of the paper [33] establish the existence, an explicit construction, and the exactness properties of the colocalization functors in contramodule categories in a generality far surpassing the needs of the present paper.
@@
The definitions of the direct and inverse images of quasi-coherent torsion sheaves and locally contraherent cosheaves of contramodules under morphisms of locally Noetherian formal schemes are worked out at the end of the paper. Specifically, we define the direct images of locally contraherent cosheaves of contramodules under suitable affineness conditions, and the inverse images of locally contraherent cosheaves of contramodules under suitable tightness and (very) flatness conditions on the morphism of formal schemes $\mathfrak{f}\colon\mathfrak{Y}\longrightarrow\mathfrak{X}$. The locality properties of the classes of tight and (very) flat morphisms of formal schemes are also established, based on an extensive discussion in the preliminaries.
@@ 2.1. Adic topological rings and their maps
Let $R$ be an adic topological ring and $I\subset R$ be an ideal of definition of $R$. Then, for every $n\geq 1$, the ideal $I^{n}\subset R$ is an ideal of definition, too. Another given finitely generated ideal $J\subset R$ is an ideal of definition if and only if there exist integers $n$, $m\geq 1$ such that $I^{n}\subset J$ and $J^{m}\subset I$. For any two ideals of definition $I$, $J\subset R$, the ideals $I+J$ and $IJ\subset R$ are also ideals of definition in $R$.
@@
We will say that an adic topological ring $R$ is adically Noetherian (cf. [32, Section 9]) if it satisfies the equivalent conditions of Lemma 2.1.2.
@@
Let $R$ and $S$ be two commutative rings with adic topologies, and let $f\colon R\longrightarrow S$ be a ring homomorphism. Then the following seven conditions are equivalent:
@@
    for every ideal of definition $I\subset R$ there exists an ideal of definition $J\subset S$ such that $J\subset Sf(I)$;
  5. (5)
    for every open ideal $I\subset R$, the ideal $Sf(I)$ is open in $S$;
  6. (6)
    the open ideals $I\subset R$ for which the ideal $Sf(I)$ is open in $S$ form a base of neighborhoods of zero in $R$;
  7. (7)
    there exists an ideal of definition $I\subset R$ such that $Sf(I)$ is an open ideal in $S$.
@@
(3) $\Longrightarrow$ (4) $\Longrightarrow$ (1) and (3) $\Longrightarrow$ (2) $\Longrightarrow$ (1) Obvious.
(4) $\Longrightarrow$ (5) Given an open ideal $I\subset R$, there exists an ideal of definition $I^{\prime}\subset R$ such that $I^{\prime}\subset I$. By (4), there exists an ideal of definition $J^{\prime}\subset S$ such that $J^{\prime}\subset Sf(I^{\prime})$. Then $J^{\prime}\subset Sf(I)$, and it follows that $Sf(I)$ is an open ideal in $S$.
(5) $\Longrightarrow$ (6) $\Longrightarrow$ (7) $\Longleftrightarrow$ (1) Obvious.
(6) $\Longrightarrow$ (4) Let $I\subset R$ be an ideal of definition. By (6), there exists an open ideal $I^{\prime}\subset R$ such that $I^{\prime}\subset I$ and the ideal $Sf(I^{\prime})$ is open in $S$. Then there exists an ideal of definition $J\subset S$ such that $J\subset Sf(I^{\prime})$. It follows that $J\subset Sf(I)$. ∎
@@ 2.2. Torsion modules
Let $R$ be a commutative ring and $s\in R$ be an element. Following the notation in [18, Section 1.1], [21, Section 2.1], etc., we denote by $R[s^{-1}]$ the ring $R$ with the element $s$ formally inverted. In other words, $R[s^{-1}]=S^{-1}R$ is the localization of $R$ at the multiplicative subset $S=\{1,s,s^{2},s^{3},\dotsc\}\subset R$.
@@
Let $I\subset R$ be an ideal. We say that an $R$-module $\mathcal{M}$ is $I$-torsion if $\mathcal{M}$ is $s$-torsion for all $s\in R$. One can easily see that it suffices to check this condition for the element $s$ ranging over any chosen set of generators of the ideal II[31, Lemma 1.2].
Other definitions of the term “$I$-torsion $R$-module” exist in the literature (see, e. g., [15, Section 3]), but the difference between them does not manifest itself for finitely generated ideals $I$, which we are really interested in. So, for a finitely generated ideal $I\subset R$, an $R$-module $\mathcal{M}$ is $I$-torsion if and only if, for every element $x\in\mathcal{M}$, there exists an integer $n\geq 1$ such that $I^{n}x=0$ in $\mathcal{M}$[31, Lemma 1.3].
@@
The assertions is obvious; see [21, Lemma 6.1(a)] for some details. ∎
@@
An $R$-module $\mathcal{M}$ is torsion if and only if, for every element $x\in\mathcal{M}$, the annihilator of $x$ in $R$ is an open ideal. In other words, an $R$-module $\mathcal{M}$ is torsion if and only if it is discrete (as a module over the topological ring $R$) in the sense of [40, Section VI.4], [3, Section 1.4], [16, Section D.5.1], [18, Section E.2], [19, Sections 1.4 and 2.1], [23, Section 2.3], [24, Section 2.4], [36, Section 7.2], [37, Section 1], etc.
@@
For any ideal $I$ in a commutative ring $R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ is closed under submodules, quotients, extensions, and infinite direct sums in $R{\operatorname{\mathsf{--Mod}}}$. In other words, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-tors}}}}$ is a localizing subcategory, or in a different terminology, a hereditary torsion class [40, Sections VI.2–3] in $R{\operatorname{\mathsf{--Mod}}}$.
@@ 2.3. Adic completions
There is a natural completion map $\lambda_{I,M}\colon M\longrightarrow\Lambda_{I}(M)$. We say that an $R$-module $M$ is $I$-adically separated if the map $\lambda_{I,M}$ is injective, and that $M$ is $I$-adically complete if $\lambda_{I,M}$ is surjective. For any $R$-module $M$, the $R$-module $\Lambda_{I}(M)$ is $I$-adically separated and complete [41, Corollary 3.6]. The latter assertion depends on the assumption that the ideal $I\subset R$ is finitely generated [41, Example 1.8]. The functor $\Lambda_{I}$ is left adjoint to the inclusion of the full subcategory of $I$-adically separated and complete $R$-modules into $R{\operatorname{\mathsf{--Mod}}}$, and the map $\lambda_{I,M}$ is the adjunction unit [21, Theorem 5.8].
@@
The converse assertion to Lemma 2.3.1(a) is not true: There exist continuous maps of adic topological rings $f\colon R\longrightarrow S$ such that the continuous ring map $\Lambda(f)$ is tight but the map $f$ is not. For a counterexample, it suffices to take $R=S=k[x]$ and $f=\mathrm{id}_{k[x]}$ as in Examples 2.1.7, with the discrete ($(0)$-adic) topology on $R$ and the indiscrete ($(1)$-adic) topology on $S$. Then one has $\Lambda(R)=R$ and $\Lambda(S)=0$; so the ring map $\mathfrak{f}=\Lambda(f)$ is tight; but the ring map $f$ is not. We will continue this discussion in Section 2.10; see Corollary 2.10.4 and Proposition 2.10.7.
Let $R$ be an adic topological ring and $\mathfrak{R}=\Lambda(R)$ be its completion. The adic completion functor $\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ or $\Lambda\colon R{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ is neither left, nor right exact [41, Example 3.20]. This fact is closely related to the fact that the category of separated and complete $R$-modules is not abelian [21, Example 2.7(1)]. To overcome these difficulties, one considers contramodule $R$-modules instead of the separated and complete $R$-modules; see the next section.
@@ 2.4. Contramodules
Let $R$ be a commutative ring and $s\in R$ be an element. An $R$-module $P$ is said to be an $s$-contramodule if $\operatorname{Hom}_{R}(R[s^{-1}],P)=0=\operatorname{Ext}_{R}^{1}(R[s^{-1}],P)$. The ring $R[s^{-1}]$ is considered as an $R$-module here. The definition above does not mention any $\operatorname{Ext}^{i}$ with $i\geq 2$ because the projective dimension of the $R$-module $R[s^{-1}]$ never exceeds 11[18, Section 1.1], [17, proof of Lemma B.7.1], [21, proof of Lemma 2.1]. A detailed discussion of this definition, which goes back to [16, Remark A.1.1], [17, Theorem B.1.1], [18, Section D.1], and [20, Section 2], can be found in [21, Sections 1–3].
Let $I\subset R$ be an ideal. An $R$-module $\mathfrak{P}$ is said to be an $I$-contramodule (or an $I$-contramodule $R$-module) if $\mathfrak{P}$ is an $s$-contramodule for every $s\in I$. It suffices to check this condition for the element $s$ ranging over any chosen set of generators of the ideal II[17, Theorem B.1.1], [20, Section 2]. A detailed discussion with two proofs can be found in [21, Theorem 5.1].
@@
This is [21, Lemma 6.1(b)]. ∎
Assume that the ideal $I\subset R$ is finitely generated. Then any $I$-contramodule $R$-module is $I$-adically complete [21, Theorem 5.6], but it need not be $I$-adically separated [38, Example 2.5], [41, Example 3.20], [15, Example 4.33], [16, Section A.1.1], [19, Section 1.5], [21, Example 2.7(1)]. Any $I$-adically separated and complete $R$-module is an $I$-contramodule [21, Lemma 5.7].
An $I$-contramodule $R$-module is said to be quotseparated [35, Section 5.5], [25, Section 1] if it is a quotient $R$-module of an $I$-adically separated and complete $R$-module. The references in the previous paragraph provide examples of quotseparated $I$-contramodule $R$-modules that are not $I$-adically separated. For an example of an $I$-contramodule $R$-module that is not quotseparated, see [20, Example 2.6], [25, Examples 1.8]. Over a Noetherian commutative ring $R$, all $I$-contramodule $R$-modules are quotseparated [17, Theorem B.1.1], [25, Corollary 3.7 and Remark 3.8]. A perhaps more detailed discussion can be found in [32, Section 1].
Now let $R$ be a commutative ring endowed with an adic topology in the sense of Section 2.1. Let $I$, $J\subset R$ be two ideals of definition of the ring $R$. Then an $R$-module $\mathfrak{P}$ is an $I$-contramodule if and only if it is a $J$-contramodule [21, Remark 5.5]. If this is the case, we will say that $\mathfrak{P}$ is a contramodule $R$-module (with respect to the given adic topology of $R$). Moreover, a contramodule $R$-module is quotseparated as an $I$-contramodule $R$-module if and only if it is quotseparated as a $J$-contramodule $R$-module; so one can speak of quotseparated contramodule $R$-modules.
@@
For any ideal $I$ in a commutative ring $R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is closed under kernels, cokernels, extensions, and infinite direct products in $R{\operatorname{\mathsf{--Mod}}}$[8, Proposition 1.1], [20, Section 2], [21, Theorem 1.2(a)]. It follows that $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is an abelian category. In fact, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is a locally $\aleph_{1}$-presentable abelian category with enough projective objects [34, Example 4.1(3)] (see also [22, Example 1.3(4)]). The identity inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is exact and preserves infinite products (hence also all limits) and $\aleph_{1}$-directed colimits, but not infinite direct sums. The infinite direct sum functors in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ are usually not exact.
@@
In the case of a finitely generated ideal $I\subset R$, the full subcategory $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ is closed under kernels, cokernels, and infinite direct products in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ and in $R{\operatorname{\mathsf{--Mod}}}$, and also under subobjects and quotient objects in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$[25, Lemma 1.2]. (However, the full subcategory of quotseparated contramodules is not closed under extensions; in fact, any $I$-contramodule $R$-module is an extension of two quotseparated $I$-contramodule $R$-modules [25, Proposition 1.6].) It follows that $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ is an abelian category. In fact, $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ is also a locally $\aleph_{1}$-presentable abelian category with enough projective objects [25, end of Section 1]. The identity inclusion functors $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}$ and $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ are exact and preserve infinite products (hence also all limits) and $\aleph_{1}$-directed colimits. The inclusion functor $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ does not preserve infinite direct sums. The infinite direct sum functors in $R{\operatorname{\mathsf{--Mod}}}_{I{\operatorname{\mathsf{-ctra}}}}^{\mathsf{qs}}$ are usually not exact (the global dimension $1$ case is one major exception [17, Remark 1.2.1]).
@@
To any complete, separated topological ring $\mathfrak{R}$ where open right ideals form a base of neighborhoods of zero, the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ of left $\mathfrak{R}$-contramodules is naturally assigned [16, Remark A.3], [17, Section 1.2], [19, Section 2.1], [34, Sections 1.2 and 5], [23, Section 2.7], [24, Sections 2.6–2.7], [36, Section 6], [37, Section 1]. The category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ comes equipped with a natural exact, faithful forgetful functor $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$, preserving infinite products (hence all limits).
@@
In the case of an adic topological ring $R$, the composition of forgetful functors $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful, and its essential image coincides with the full subcategory of quotseparated contramodule $R$-modules [25, Proposition 1.5]. So there is a natural equivalence (in fact, isomorphism) of categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}\simeq R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}$.